OpenStudy (anonymous):

consider the equations f(x)=-6x-1 and g(x)=4x^2 select the solution for (f+g)(x)

OpenStudy (anonymous):

f(x) =-6-1=-7 therefore. (fg)(x)=-7(4x^2)=-28x^2.

OpenStudy (marissalovescats):

It's asking for (f+g)(x) So this means you do f(x)+g(x)

OpenStudy (marissalovescats):

AKA just combining like terms.

OpenStudy (marissalovescats):

 There are no like terms to combine so you're just combining it.